# Blockchain with proof of quantum work

### v1

@@ Abstract
We propose a blockchain architecture in which mining requires a quantum computer. The consensus mechanism is based on proof of quantum work, a quantum-enhanced alternative to traditional proof of work that leverages quantum supremacy to make mining intractable for classical computers. We have refined the blockchain framework to incorporate the probabilistic nature of quantum mechanics, ensuring stability against sampling errors and hardware inaccuracies. To validate our approach, we implemented a prototype blockchain on four D-WaveTM quantum annealing processors geographically distributed within North America, demonstrating stable operation across hundreds of thousands of quantum hashing operations. Our experimental protocol follows the same approach used in the recent demonstration of quantum supremacy [1], ensuring that classical computers cannot efficiently perform the same computation task. By replacing classical machines with quantum systems for mining, it is possible to significantly reduce the energy consumption and environmental impact traditionally associated with blockchain mining. Beyond serving as a proof of concept for a meaningful application of quantum computing, this work highlights the potential for other near-term quantum computing applications using existing technology.
@@ I Introduction
Unlike previous quantum blockchain proposals [17, 18, 19, 20, 21, 22, 23, 24, 25, 26], which often assume access to fault-tolerant quantum computing or quantum teleportation, our approach integrates both classical and quantum computation to leverage quantum supremacy with existing and near-term QPUs. Currently, Ref. [1] is the only demonstration of quantum supremacy not based on random-circuit sampling [27, 28], and thus the only quantum supremacy demonstration relevant to this work (see Appendix A). Using a small-scale prototype, we demonstrate this blockchain’s functionality on four generally-accessible D-Wave AdvantageTM and Advantage2TM quantum processing units (QPUs) at different geographic locations across North America. By carefully accounting for the probabilistic nature of quantum computation and potential errors within Bitcoin’s architecture, we achieve stable operation across thousands of mining processes distributed among hundreds of stakeholders. This marks the first successful demonstration of an important real-world application that directly leverages quantum supremacy, operating seamlessly across an internationally distributed network of quantum computers.
@@ II Classical Blockchain Architecture
Before introducing any quantum concepts, first described how standard blockchain technology works, using Bitcoin’s architecture as a prime example. At the heart of Bitcoin is its distributed ledger, a decentralized database that records all transactions. This ledger is maintained across a network of computers, ensuring both transparency and security. Each node on the network stores a complete copy of the blockchain, and updates are synchronized through consensus mechanisms. These mechanisms ensure that only valid transactions are added to the blockchain, making it resistant to tampering or fraud.
Figure 1: Illustration of Bitcoin blocks.
@@
A situation known as a “fork” occurs when two miners solve a block at nearly the same time, resulting in the blockchain temporarily splitting into two competing branches. Both branches propagate through the network, and nodes may accept one branch over the other depending on which block they receive first. The fork is resolved when miners adding new blocks to one of the branches make it longer than the other. Bitcoin follows the “strongest chain rule”, meaning the branch with the highest total computational work (often called Chainwork) is recognized as the valid chain, while the shorter branch is discarded. This process ensures network consensus but can result in transaction delays and unrewarded work. It is therefore essential to understand how the strongest chain is determined.
@@
Chainwork is calculated by the stakeholders for all existing chains and the chain with largest Chainwork is selected for adding the next block and the rest are discarded. The miners responsible for the discarded blocks do not receive rewards, while the miner who successfully extends the longest chain is rewarded with both block rewards and transaction fees from the new block. Transactions from the discarded branch are returned to the memory pool (mempool), where they await inclusion in a future block. This mechanism ensures that consensus is maintained across the network, although it may introduce temporary delays in transaction confirmation. On a strongest chain, blocks closer to the genesis block are increasingly likely to remain unchanged as broadcasts continue. If we define a small cut-off in probability, we might say blocks become (for practical purposes) immutable, and define the transaction delay accordingly. In the case of Bitcoin it is only a few blocks, assuming good network synchronization. Miners are incentivized to maintain synchronization, to minimize the risk of pursuing wasteful computations, so that forks are merely a rare inconvenience and ${\cal D}$ of 2 or 3 is typical. Forks may also happen due to software updates and change of rules, which may lead to temporary (soft-fork) or permanent (hard-fork) splitting of the blockchain.
We highlight three features required for viability of a blockchain: 1. The number of blocks in the strongest chain relative to the number of honestly-mined blocks broadcast should be a large finite faction (the efficiency). 2. A stakeholder should have confidence that that blocks buried to some small finite depth on the strongest chain are immutable with high probability. The ratio depth/efficiency dictates the expected number of honestly-mined blocks that have to be broadcast before the most recently broadcast transactions can be considered reliably processed. 3. Consensus mechanisms should protect against disruptions, and prevent rewards from being obtained (on average) by incomplete or fraudulent work. These are necessary requirements for guaranteeing confidence, and maintaining an incentive to mine according to protocol rules.
@@ IV Quantum Hashing
    where $\theta(x)$ is the heaviside step-function and $W_{0\alpha}$ is a threshold selected depending on the definition of $W_{\alpha}$. Hash is a sequence of the binaries:
@@
The quantum operation must be sufficiently complex to satisfy the seventh requirement, ensuring that the hash cannot be spoofed by classical computers with feasible resources. Such complexity can be achieved through a continuous-time evolution, sequence of quantum gate operations, or a combination of digital and analog operations. Information from the resulting density matrix is extracted through measurements and summarized in a vector of witnesses $W(\rho_{\mathcal{M}})$. It should be noted that some classical properties, such as collision resistance and the avalanche effect, do not necessarily carry over automatically. For example, if quantum evolution maps all $\Theta$ to the same hash value, these properties would be lost. One should therefore be careful in the design of quantum evolution and the witness vector. The definition of witnesses and the quality of the resulting hash ultimately depend on the quantum hardware’s capability to execute complex quantum evolutions and its measurement capabilities, as we discuss next.
@@ IV.1 Single-Basis Measurement
IV.2 Multi-Basis Measurement
As an illustrative example, we consider using Bell’s inequality violations for pairs of qubits as entanglement witnesses. In an appropriate experimental setup, a loophole-free violation of Bell’s inequality indicates that the experimental results cannot be explained by any local hidden-variable theory. However, this is not our objective. Instead, we aim to use the inequality as an entanglement witness. Specifically, we employ an experimentally feasible variation of Bell’s inequality, known as the Clauser-Horne-Shimony-Holt (CHSH) inequality. Additional details are provided in Appendix 35.
Let us consider a system in which measurements can be performed in two Pauli bases, $x$ and $z$. For any two qubits, $i$ and $j$, we define a modified Bell operator as (see also Appendix B):
$B_{ij}=\sqrt{2}(\sigma_{i}^{x}\sigma_{j}^{x}+\sigma_{i}^{z}\sigma_{j}^{z}).$ (11)
Based on purely statistical argument, it is shown that for any local hidden variable theory, the expectation value of the Bell operator satisfies
$|\langle B_{ij}\rangle|\leq 2\;,$ (12)
which is known as CHSH inequality. When Pauli matrices are replaced by classical spin components, we get an even smaller bound, $|\langle B_{ij}\rangle|\leq\sqrt{2}$ (see Appendix B). In contrast, quantum mechanics allows $|\langle B_{ij}\rangle|$ to reach values as large as $2\sqrt{2}$. This is possible when $|\langle\sigma_{i}^{x}\sigma_{j}^{x}\rangle|=|\langle\sigma_{i}^{z}\sigma_{j}^{z}\rangle|=1$, which is classically counterintuitive as it means the spins are oriented both in $x$ and $z$ directions at the same time. Quantum mechanically it is possible because of quantum contextuality, i.e., dependence of the outcome on the measurement setup. It is clear that violation of CHSH inequality is a sufficient but not necessary condition for quantum mechanics, and more explicitly for entanglement between the qubit pair. Therefore, CHSH inequality is an entanglement witness.
CHSH inequality can also hold in a multi-qubit system and therefore can be used as a basis for hash generation. Let us define $W_{\alpha}=|\langle B_{ij}\rangle|$ in Eq. (6) with a threshold value $W_{0\alpha}=2$. Here, $\alpha$ represents a pair of qubits in the multi-qubit system. The hash bit ${\cal H}_{\alpha}$ is then equal to one if the CHSH inequality for the corresponding pair is violated, and zero otherwise. It has been shown that in a multi-qubits system the CHSH inequality must satisfy a trade-off relation [45] (see Appendix B):
$\sum_{\alpha\in\{{\rm all\ pairs}\}}W_{\alpha}^{2}\ \leq 2N_{Q}(N_{Q}-1)$ (13)
IV.3 Shadow Tomography
$\langle O\rangle\approx{1\over N}\sum_{i}{\rm Tr}(O\hat{\rho}_{i})$ (14)
where $M$ is the shadow tomography measurement channel, defined as:
${\rm Tr}(O_{\alpha}\rho)>\kappa>{\rm Tr}(O_{\alpha}\tilde{\rho})$ (18)
@@ V.1 Probabilistic Validation
represents a dimensionless distance from the corresponding hyperplane. As expected, $\varepsilon_{\alpha}\to 0$ for $d_{\alpha}\gg 1$ and $\varepsilon_{\alpha}\to 0.5$ for $d_{\alpha}\to 0$. It is more convenient to work with confidence values defined as ${\cal P}_{\alpha}\equiv 1-\varepsilon_{\alpha}$, instead of error rates. Assuming that errors in witnesses are uncorrelated (this assumption does not affect the applicability of this method as a heuristic), the miner’s confidence in the generated ${\cal N_{H}}$-bit hash value is given by
@@ V.2 Redefining Chainwork
Scenario (1) would lead to a soft validation fork. The miner’s proposed branch becomes obsolete as a new branch mined by the dominant computing power quickly replaces it. The isolated miner rejoins the strongest chain with high probability after two block broadcasts. This soft fork is similar to those caused by network desynchronization in Bitcoin and is not problematic.
Scenario (2), on the other hand, can cause a hard validation fork. A stakeholder who fails to validate a block is left waiting for a new fork that may emerge slowly but is unlikely to be considered the strongest chain by others. Since the stakeholder cannot join a chain with invalid transactions, they become permanently disconnected from the main chain. If this issue persists, stakeholders will eventually split into hard forks, causing the blockchain to stop functioning. To prevent this, the definition of the strongest chain must allow for the inclusion of unverified blocks.
@@
If the work remains constant, then one can assign ±$\pm 1$ to valid/invalid blocks, we call this basic +/- Chainwork. In the case of a tie, the first path evaluated is preferred, following the same rule as in Bitcoin. This adjustment effectively resolves the issue by turning a hard validation fork to a soft validation fork with a lag time. If a user falls behind by two blocks, the majority compute power of miners will produce a third block, causing the strongest chain to advance by three blocks in agreement with the majority.
@@ VI Experimental Implementation
Experimental realization of a quantum blockchain requires access to a quantum computer capable of solving complex problems where quantum supremacy has been demonstrated or is potentially achievable. The generated distribution must exhibit sufficient structure to allow observables with nontrivial expectation values that can be measured reproducibly across multiple QPUs. Therefore, techniques such as boson sampling or random circuit sampling, which produce nearly uniform random outcomes, are not suitable for this application (see Appendix A). Currently, Ref. [1] is the only experimental demonstration of quantum supremacy that meets these requirements. Therefore, we base our experimental demonstration on the same protocol. Since Ref. [1] has already established agreement with quantum mechanics and the infeasibility of classical simulation, our primary focus here is to demonstrate that multiple QPUs can generate and validate hashes, ensuring the blockchain operates reliably without persistent branching or network fragmentation.
Figure 3: Operation of an example blockchain with $100$ miners using four publicly accessible Advantage and Advantage2 QPUs. The mining difficulty was set to $\mathcal{N}_{\rm zeros}=28$ with basic +/-1 Chainwork defined in (26). Orange blocks show soft (resolved) forks, these blocks are included in no strongest chains (have been rejected by all miners). Blue blocks are immutable, included in all strongest chains (agreed upon by all miners). Gray and black blocks are in contention: common to some (but not all) strongest chains. Only black blocks are candidates for maximum Chainwork (are being mined) and have the potential for further branching. Blocks are labeled by the order of broadcast with the genesis block on the left and most recently mined block on the right. Either of the paths terminating in black nodes might be extended, but the orange path is no longer a candidate for the strongest chain. Blocks 10, 11, 12 and 13 contain delayed transactions, that require further mining (fork resolution) in order to be confirmed. A majority of miners believe node 13 has maximum Chainwork, and in this experiment all miners consolidate on this branch as the chain develops (see Figure 4).
@@
Figure 5: Mining efficiency, i.e., the fraction of blocks joining the strongest chain, from 5-10 chains each of length 512-2048. The Chainwork uses either the simple +/-1 weighting, or the confidence-based weighting with $\mathcal{N}_{\rm max}=1,2$ or $3$ with a constant $\delta W_{\alpha}$ extracted from experimental data. The mining process is accelerated without impact on statistical outcomes, and witnesses are resampled to accelerate analysis. Further experimental details and results are contained in Appendix D. Confidence-weighted strongest chains are more efficient.
@@
The feature vector is constructed to include only pairwise nearest-neighbor correlations: $V=[...,\langle\sigma^{z}_{i}\sigma^{z}_{j}\rangle,...]^{T}$. As demonstrated in [1], reproducing $V$ within the accuracy of QPUs is computationally infeasible for classical computers. The witness vector $W$ is calculated using random projection by a matrix $G$ with matrix elements selected from a normal distribution $N(0,I)$; elements are pseudo-random functions of ${\mathcal{M}}$. As a miner, we run the quantum algorithm multiple times, each time with a different nonce until we find a hash with $\mathcal{N}_{\rm zeros}$ bits equal to zero (more specifically for confidence-based Chainwork, a hash satisfying the confidence threshold). The miner then adds the block to the blockchain with the nonce that generated ${\cal H}=0$. The validator then goes through the same calculations, i.e., generates the classical hash and thereby the QPU parameters and the quantum hash and checks if all bits are zeros.
@@
Fig. 3 presents the initial-stage operation of a blockchain using four publicly accessible D-Wave Advantage and Advantage2 QPUs, involving 100 miners with $\mathcal{N}_{\rm zeros}=28$. The visualization captures the dynamic evolution of the blockchain, highlighting block statuses and mining activity. Blocks on the (majority) strongest chain are arranged chronologically from left to right, with the genesis block on the far left and the most recently mined block on the far right. Different block colors indicate their status within the consensus process: Blue blocks (immutable) are included in all strongest chains and are universally agreed upon. Orange blocks (soft forks) have been rejected by all miners and are not included in any of the strongest chains. Gray and black blocks appear in some but not all strongest chains, reflecting temporary disagreement among miners. Only black blocks are actively being mined, meaning they have the potential for further branching. This illustrates how consensus emerges dynamically as blocks are proposed and either accepted or rejected by the network.
@@
In Fig. 4, the blockchain is shown after 219 block broadcasts, with the genesis block repositioned centrally for clarity. By this stage, approximately 70% of proposed blocks have become immutable (blue), aligning well with theoretical expectations based on the resampling methodology (see Appendix D.3). The remaining blocks continue to be evaluated, with only those on the strongest chains persisting in the blockchain ledger. This simulation demonstrates the robustness of the blockchain framework, showcasing its ability to maintain stability and achieve miner consensus despite the probabilistic nature of quantum computation. We are able to use statistical properties of the Chainwork described in Section D.2 to eliminate failed mining experiments, and create the blockchain with only $100\times 219$ total QPU programmings.
@@
Fig. 5 illustrates the mining efficiency across multiple blockchain operations. Mining efficiency is defined as the fraction of block broadcasts that successfully enter the strongest chain in the limit of many miners. For each curve the data is taken from 5 to 10 chains, each having between 512 and 2048 blocks. In Appendix D.1 we describe mining efficiency and transaction delay in greater detail, presenting additional data. The analysis compares different Chainwork definitions and post-processing methods. For the simple ±$\pm 1$ weighting method, which assigns equal weights to all block broadcasts, mining efficiency drops to close to 50% as $\mathcal{N}_{\rm zeros}\to 60$. The confidence-based weighting technique performs better, as expected, approaching 75% efficiency for $\mathcal{N}_{\rm zeros}=60$. Confidence-based Chainwork allows higher efficiency than basic +/-1 Chainwork, also with accounting of small differences in the mining rate as outlined in Appendix D.2. To calculate confidence values we used a constant $\delta W_{\alpha}$ extracted from experimental data, as detailed in Appendix D.

### v2

@@ Abstract
We propose a blockchain architecture in which mining requires a quantum computer. The consensus mechanism is based on proof of quantum work, a quantum-enhanced alternative to traditional proof of work that leverages quantum supremacy to make mining intractable for classical computers. We have refined the blockchain framework to incorporate the probabilistic nature of quantum mechanics, ensuring stability against sampling errors and hardware inaccuracies. To validate our approach, we implemented a prototype blockchain on four D-WaveTM quantum annealing processors geographically distributed within North America, demonstrating stable operation across hundreds of thousands of quantum hashing operations. Our experimental protocol follows the same approach used in the recent demonstration of quantum supremacy [1], ensuring that classical computers cannot efficiently perform the same computation task. By replacing classical machines with quantum systems for mining, it is possible to significantly reduce the energy consumption and environmental impact traditionally associated with blockchain mining while providing a quantum-safe layer of security. Beyond serving as a proof of concept for a meaningful application of quantum computing, this work highlights the potential for other near-term quantum computing applications using existing technology.
@@ I Introduction
Unlike some previous quantum blockchain proposals [17, 18, 19, 20, 21, 22, 23, 24, 25, 26], which often assume access to fault-tolerant quantum computing or quantum teleportation, our approach integrates both classical and quantum computation to leverage quantum supremacy with existing and near-term QPUs. Our proposal is also different from blockchains and hashes that have been proposed to exploit the potential of near-term boson-sampling [27, 28, 29]. We introduce a hierarchy of practical witness methods, new consensus mechanisms robust to noise, and implement a proof-of-concept experiment using a distributed network of quantum computers.
Although claims of supremacy have been made in Boson sampling, these have been challenged [30, 31]. Currently, Ref. [1] is the only demonstration of quantum supremacy not based on random-circuit sampling [32, 33], and thus the only demonstration of quantum supremacy relevant to this work (see Appendix A). The unitary dynamics used for hashing remains intractable despite recent developments in algorithmic methods [34, 35, 36].
Using a small-scale prototype, we demonstrate this blockchain’s functionality on four generally-accessible D-Wave AdvantageTM and Advantage2TM quantum processing units (QPUs) at different geographic locations across North America. By carefully accounting for the probabilistic nature of quantum computation and potential errors within Bitcoin’s architecture, we achieve stable operation across thousands of mining processes distributed among hundreds of stakeholders. This marks the first successful demonstration of an important real-world application that directly leverages quantum supremacy, operating seamlessly across an internationally distributed network of quantum computers.
@@ II Classical Blockchain Architecture
Before introducing any quantum concepts, we first describe how standard blockchain technology works, using Bitcoin’s architecture as a prime example. At the heart of Bitcoin is its distributed ledger, a decentralized database that records all transactions. This ledger is maintained across a network of computers, ensuring both transparency and security. Each node on the network stores a complete copy of the blockchain, and updates are synchronized through consensus mechanisms. These mechanisms ensure that only valid transactions are added to the blockchain, making it resistant to tampering or fraud.
Figure 1: Illustration of Bitcoin blocks. The Bitcoin block structure is sufficient for our proof-of-concept implementation; modifications may be useful for mitigation of weaknesses as discussed in Appendix C.
@@
A situation known as a “fork” occurs when two miners solve a block at nearly the same time, resulting in the blockchain temporarily splitting into two competing branches. Both branches propagate through the network, and nodes may accept one branch over the other depending on which block they receive first. The fork is resolved when miners adding new blocks to one of the branches make it longer than the other. Bitcoin follows the “strongest chain rule”, meaning the branch with the highest total computational work (often called Chainwork) is recognized as the valid chain, while the shorter branch is discarded. This process ensures network consensus, but can result in transaction delays and unrewarded work. It is therefore essential to understand how the strongest chain is determined.
@@
Chainwork is calculated by the stakeholders for all existing chains and the chain with largest Chainwork is selected for adding the next block and the rest are discarded. The miners responsible for the discarded blocks do not receive rewards, while the miner who successfully extends the longest chain is rewarded with both block rewards and transaction fees from the new block. Transactions from the discarded branch are returned to the memory pool (mempool), where they await inclusion in a future block. This mechanism ensures that consensus is maintained across the network, although it may introduce temporary delays in transaction confirmation. On a strongest chain, blocks closer to the genesis block are increasingly likely to remain unchanged as broadcasts continue. If we define a small cut-off in probability, we might say blocks become (for practical purposes) immutable, and define the transaction delay accordingly. In the case of Bitcoin it is only a few blocks, assuming good network synchronization. Miners are incentivized to maintain synchronization, to minimize the risk of pursuing wasteful computations, so that forks are merely a rare inconvenience and a delay of no more than 2 blocks is typical. Forks may also happen due to software updates and change of rules, which may lead to temporary (soft-fork) or permanent (hard-fork) splitting of the blockchain.
We highlight three features required for viability of a blockchain: 1. The number of blocks in the strongest chain relative to the number of honestly-mined blocks broadcast should be a large finite fraction (the efficiency). 2. A stakeholder should have confidence that that blocks buried to some small finite depth on the strongest chain are immutable with high probability. The ratio depth/efficiency dictates the expected number of honestly-mined blocks that have to be broadcast before the most recently broadcast transactions can be considered reliably processed. 3. Consensus mechanisms should protect against disruptions, and prevent rewards from being obtained (on average) by incomplete or fraudulent work. These are necessary requirements for guaranteeing confidence, and maintaining an incentive to mine according to protocol rules.
@@ IV Quantum Hashing
    where $\theta(x)$ is the heaviside step-function and $W_{0\alpha}$ is a threshold selected depending on the definition of $W_{\alpha}$. The hash is a sequence of the binaries:
@@
The quantum operation must be sufficiently complex to satisfy the seventh requirement, ensuring that the hash cannot be spoofed by classical computers with feasible resources. Such complexity can be achieved through a continuous-time evolution, sequence of quantum gate operations, or combination of digital and analog operations. Information from the resulting density matrix is extracted through measurements and summarized in a vector of witnesses $W(\rho_{\mathcal{M}})$. It should be noted that some classical properties, such as collision resistance and the avalanche effect, do not necessarily carry over automatically. For example, if quantum evolution maps all $\Theta$ to the same hash value, these properties would be lost. One should therefore be careful in the design of the quantum evolution and the witness vector. The definition of witnesses and the quality of the resulting hash ultimately depend on the quantum hardware’s capability to execute complex quantum evolutions and its measurement capabilities, as we discuss next.
@@
Multi-basis measurements and shadow tomography allow additional near-term flexibility in the hash definition. In particular we note that entanglement witnesses might be used, with thresholds in/at the entangled regime. Since classical probabilistic models do not allow for beyond-threshold statistics, this makes the experimental outcome difficult to simulate by any classical models. These methods are described in detail in Appendix B.
@@ V.1 Probabilistic Validation
We now develop a model of probabilistic validation assuming independent and Gaussian distributed witness statistics, which is consistent with witnesses being determined by a sum of many weakly correlated statistics, and proves to be a good approximation in experiments as shown later. We introduce a confidence measure that quantifies the accuracy of hash generation and validation relative to an ideal, error-free scenario. The proposed confidence measure can serve as a principled heuristic for practical applications, independent of the underlying assumptions
@@
represents a dimensionless distance from the corresponding hyperplane. As expected, $\varepsilon_{\alpha}\to 0$ for $d_{\alpha}\gg 1$ and $\varepsilon_{\alpha}\to 0.5$ for $d_{\alpha}\to 0$. It is more convenient to work with confidence values defined as ${\cal P}_{\alpha}\equiv 1-\varepsilon_{\alpha}$, instead of error rates. Assuming that errors in witnesses are uncorrelated, the miner’s confidence in the generated ${\cal N_{H}}$-bit hash value is given by
@@ V.2 Redefining Chainwork
Scenario (1) would lead to a soft validation fork. The miner’s proposed branch becomes obsolete as a new branch mined by the dominant computing power quickly replaces it. The isolated miner rejoins the strongest chain with high probability after two block broadcasts. This soft fork is similar to those caused by network desynchronization in Bitcoin and is not problematic other than it results in some unrewarded work that must be accounted for in mining rewards.
Scenario (2), on the other hand, can cause a hard validation fork. Since the stakeholder cannot join a chain with invalid transactions, they become permanently disconnected from the main chain. A stakeholder who fails to validate a block is left waiting for a new fork that may emerge slowly (or not at all), and regardless is unlikely to be considered the strongest chain by others. If this issue persists, stakeholders will eventually split into hard forks, causing the blockchain to stop functioning. To prevent this, the definition of the strongest chain must allow for the inclusion of unverified blocks.
@@
If the work remains constant, then one can assign ±$\pm 1$ to valid/invalid blocks, we call this basic +/- Chainwork. In the case of a tie, the first path evaluated is preferred, following the same rule as in Bitcoin. This adjustment effectively resolves the issue by turning a hard validation fork to a soft validation fork that is resolved subject to a delay. If a user falls behind by two blocks, the majority compute power of miners will produce a third block, causing the strongest chain to advance by three blocks in agreement with the majority.
@@ VI Experimental Implementation
Experimental realization of a quantum blockchain requires access to a quantum computer capable of solving complex problems where quantum supremacy has been demonstrated or is potentially achievable. The generated distribution must exhibit sufficient structure to allow observables with nontrivial expectation values that can be measured reproducibly across multiple QPUs. Therefore, techniques such as boson sampling or random circuit sampling which cannot be formulated as efficient decision problems, are not suitable for this application (see Appendix A). Coarse-grained boson sampling has been proposed as a means to implement blockchains [28, 29], but to our knowledge there is no experimental implementation (see Appendix A.1). Currently, Ref. [1] is the only experimental demonstration of quantum supremacy that meets the requirement of efficient cross-validation. Therefore, we base our experimental demonstration on the same protocol. Since Ref. [1] has already established agreement with quantum mechanics and the infeasibility of classical simulation, our primary focus here is to demonstrate that multiple QPUs can generate and validate hashes, ensuring the blockchain operates reliably with high efficiency and small transaction delays.
Figure 3: Operation of an example blockchain with $50$ miners using four publicly accessible Advantage and Advantage2 QPUs. The mining difficulty was set to $\mathcal{N}_{\rm zeros}=32$ with basic +/-1 Chainwork defined in (18). Orange blocks show soft (resolved) forks, these blocks are included in no strongest chains (have been rejected by all miners). Blue blocks are immutable, included in all strongest chains (agreed upon by all miners). Gray and black blocks are in contention: common to some (but not all) strongest chains. Only black blocks are candidates for maximum Chainwork (are being mined) and have the potential for further branching. Blocks are labeled by the order of broadcast with the genesis block on the left and most recently mined block on the right. Either of the paths terminating in black nodes might be extended, but the orange path is no longer a candidate for the strongest chain. Blocks 10, 11, 12 and 13 contain delayed transactions, that require further mining (fork resolution) in order to be confirmed. A majority of miners believe node 13 has maximum Chainwork, and in this experiment all miners consolidate on this branch as the chain develops (see Figure 4).
@@
Figure 5: Mining efficiency, i.e., the fraction of blocks joining the strongest chain. The Chainwork uses either the simple +/-1 weighting, or the confidence-based weighting with $\mathcal{N}_{\rm max}=1,$ $2$ or $3$ with a constant $\delta W_{\alpha}$ extracted from experimental data. Confidence-weighted strongest chains are more efficient. The mining process is accelerated without impact on statistical outcomes as described in Appendix  D.3. Results are obtained with resampling from experimentally-parameterized witness distributions to accelerate analysis, each point an average on 16 chains of length 512-1024. Resampling statistics are in agreement with full-experiment efficiencies reported in Figures 4 and 10. Further experimental details and statistics are contained in Appendix D.
@@
The feature vector is constructed to include only pairwise nearest-neighbor correlations: $V=[...,\langle\sigma^{z}_{i}\sigma^{z}_{j}\rangle,...]^{T}$. As demonstrated in [1], reproducing $V$ within the accuracy of QPUs is computationally infeasible for classical computers. The witness vector $W$ is calculated using random projection by a matrix $G$ with matrix elements selected from a normal distribution $N(0,I)$; elements are pseudo-random functions of ${\mathcal{M}}$. Thresholds ($W_{0}$) for the witnesses are set as zero. As a miner, we run the quantum algorithm multiple times, each time with a different nonce until we find a hash with $\mathcal{N}_{\rm zeros}$ bits equal to zero (more specifically for confidence-based Chainwork, a hash satisfying the confidence threshold). The miner then adds the block to the blockchain with the nonce that generated ${\cal H}=0$. The validator then goes through the same calculations, i.e., generates the classical hash and thereby the QPU parameters and the quantum hash and checks if all bits are zeros.
@@
Fig. 3 presents the initial-stage operation of a blockchain using four publicly accessible D-Wave Advantage and Advantage2 QPUs, involving 50 miners with $\mathcal{N}_{\rm zeros}=32$. The visualization captures the dynamic evolution of the blockchain, highlighting block statuses and mining activity. Blocks on the (majority) strongest chain are arranged chronologically from left to right, with the genesis block on the far left and the most recently mined block on the far right. Different block colors indicate their status within the consensus process: Blue blocks (immutable) are included in all strongest chains and are universally agreed upon. Orange blocks (soft forks) have been rejected by all miners and are not included in any of the strongest chains. Gray and black blocks appear in some but not all strongest chains, reflecting temporary disagreement among miners. Only black blocks are actively being mined, meaning they have the potential for further branching. This illustrates how consensus emerges dynamically as blocks are proposed and either accepted or rejected by the network.
@@
In Fig. 4, the blockchain is shown after 219 block broadcasts, with the genesis block repositioned centrally for clarity. By this stage, approximately 70% of proposed blocks have become immutable (blue), aligning well with theoretical expectations based on the bootstrapping methodology (see Appendix D.3). The remaining blocks continue to be evaluated, with only those on the strongest chains persisting in the blockchain ledger. This simulation demonstrates the robustness of the blockchain framework, showcasing its ability to maintain stability and achieve miner consensus despite the probabilistic nature of quantum computation. We are able to use statistical properties of the Chainwork described in Section D.2 to eliminate failed mining experiments, and create the blockchain with only $50\times 219$ total QPU programmings. Blockchains with more miners, and over $1000$ block broadcasts, are shown in Appendices.
@@
Fig. 5 illustrates the mining efficiency across multiple blockchain operations. Mining efficiency is defined as the fraction of block broadcasts, excluding the genesis block, that successfully enter the strongest chain in the limit of many miners. For each curve the data is taken from 16 chains created with bootstrapping of witness distributions, each having between 512 and 1024 block broadcasts. The bootstrapping method is in agreement with results without bootstrapping reported in other figures. In Appendix D.1 we describe mining efficiency and transaction delay in greater detail, presenting additional data. The analysis compares different Chainwork definitions and post-processing methods. For the basic ±$\pm 1$ weighting method, which assigns equal weights to all block broadcasts, mining efficiency drops to close to 50% as $\mathcal{N}_{\rm zeros}\to 64$. The confidence-based weighting technique performs better, as expected, approaching 75% efficiency for $\mathcal{N}_{\rm zeros}=64$. Confidence-based Chainwork allows higher efficiency than basic +/-1 Chainwork, also with accounting of small differences in the mining rate as outlined in Appendix D.2. To calculate confidence values we used a constant $\delta W_{\alpha}$ extracted from experimental data, as detailed in Appendix D.